# CrashEventLLM: Predicting System Crashes with Large Language Models

### v1

@@ I Introduction
Recent advancements in computing technology have caused a significant rise in our dependence on computing systems across a wide range of vital services. From banking operations to utility management, a plethora of everyday functions now rely on a diverse array of computing systems. The critical importance of minimizing system failures or downtime cannot be overstated, as it is often essential for societal function. System failures typically originate from system errors, where each critical error has the capacity to prompt system shutdowns or reboots in an effort to resolve the issues [1]. Such errors can lead to considerable financial losses and inflict significant damage on crucial information technology (IT) infrastructure. For example, within the Windows environment, users often encounter the disruptive ”blue screen of death” (BSOD), hampering device functionality [1]. In the event of a system crash in the Windows operating system (OS), a crash log is generated to aid in the analysis of the crash triggers. Critical crashes are identified and logged as event 41. This event signifies that some unforeseen activity hindered the proper shutdown of Windows. Such shutdowns may result from power supply interruptions or stop errors. System crash logs serve as a comprehensive record of an IT system’s operational activities and events, offering detailed insights. These logs are important in two folds: fault detection and failure prediction. Fault prediction focuses on swiftly identifying the indicators of the critical failures as they arise, often employing anomaly detection methods. On the contrary, failure prediction adopts a proactive approach, aiming to issue early alerts regarding potential failures.
@@
Several research work addressed the failure prediction problem based on machine learning and deep learning techniques [2, 3]. Recent growth of large language models (LLMs) [4, 5, 6] in the area of real world event-prediction and text-reasoning tasks have demonstrated interesting results [7]. In particular, due to being pre-trained on text, logs, and code corpora, LLMs have also shown promising effectiveness in software engineering tasks, such as software comprehension [8, 9], code generation and bug resolving [10, 11, 12, 13], log analysis [14, 15, 16, 17, 18, 19], and failure mode classification [20]. Therefore, it becomes interesting to investigate their capability in system crash prediction. In this paper, we investigate their capabilities in predicting and reasoning about the system crash events given the past. Specifically, we model system’s sequences of crash timestamps and predict their future crashes along with the possible cause of failure. Our hypothesis is that LLMs are potentially useful for advancing solutions to this problem because event sequences are often accompanied with rich text information which LLMs can handle.
@@ IV Limitations and Future Directions
For future research in system crash prediction, we propose three directions: First, fine-tuning more advanced LLMs with specific data and conducting large-scale studies could unlock new arenas for system crash prediction using crash logs and LLMs. Second, expanding the dataset size may enhance performance. Third, while we evaluated predicted outputs using ROUGE scores in this paper, advancements in LLMs warrant exploring additional performance metrics for evaluation purposes.

### v2

@@ I Introduction
Recent advancements in computing technology have caused a significant rise in our dependence on computing systems across a wide range of vital services. From banking operations, utility management, airlines, to information technology infrastructure, a plethora of everyday functions now rely on a diverse array of computing systems. The critical importance of minimizing system failures or downtime cannot be overstated, as it is often essential for societal function. System failures or crashes typically originate from system errors, where each critical error has the capacity to prompt system shutdowns or reboots in an effort to resolve the issues [1]. Such errors can lead to considerable financial losses and inflict significant damage on crucial information technology (IT) infrastructure. Within the Windows environment, users often encounter the disruptive ”blue screen of death” (BSOD), hampering device functionality [1]. Recent such example is from crowdstrike incident [2], where roughly 8.5 million systems crashed and were unable to properly restart in leading to the largest outage in the history of information technology. Typically, in the event of a system crash in the Windows operating system (OS), a crash log is generated to aid in the analysis of the crash triggers. Critical crashes are identified and logged as event 41. This event signifies that some unforeseen activity hindered the proper shutdown of Windows. Such shutdowns may result from power supply interruptions or stop errors. System crash logs serve as a comprehensive record of an IT system’s operational activities and events, offering detailed insights. These logs are important in two folds: fault detection and failure prediction. Fault prediction focuses on swiftly identifying the indicators of the critical failures as they arise, often employing anomaly detection methods. On the contrary, failure prediction adopts a proactive approach, aiming to issue early alerts regarding potential failures.
@@
Several research work addressed the failure prediction problem based on machine learning and deep learning techniques [3, 4]. A related study utilized telemetry data for system failure prediction through a concurrent work [5]. Recent growth of large language models (LLMs) [6, 7, 8] in the area of real world event-prediction and text-reasoning tasks have demonstrated interesting results [9]. In particular, due to being pre-trained on text, logs, and code corpora, LLMs have also shown promising effectiveness in software engineering tasks, such as software comprehension [10, 11], code generation and bug resolving [12, 13, 14, 15], log analysis [16, 17, 18, 19, 20, 21], and failure mode classification [22]. Therefore, it becomes interesting to investigate their capability in system crash prediction. In this paper, we investigate their capabilities in predicting and reasoning about the system crash events given the past. Specifically, we model system’s sequences of crash timestamps and predict their future crashes along with the possible cause of failure. Our hypothesis is that LLMs are potentially useful for advancing solutions to this problem because event sequences are often accompanied with rich text information which LLMs can handle.
@@ IV Limitations and Future Directions
For future research in system crash prediction, we propose three directions: First, fine-tuning more advanced LLMs with specific data and conducting large-scale studies could unlock new arenas for system crash prediction using crash logs and LLMs. Second, expanding the dataset size may enhance performance. Specifically, Mudgal et al. [5] utilized system telemetry data to detect system failures. Similarly, integrating telemetry data including system characteristics and software updates with system crash information could enhance the identification of correlations between telemetry data and system crashes, leading to more accurate predictions of system failures or crashes. Third, while we evaluated predicted outputs using ROUGE scores in this paper, advancements in LLMs warrant exploring additional performance metrics for evaluation purposes.